The combustion of methane gas is described by the equation CH4+2O2→CO2+2H2O.
What volume of oxygen is needed at STP to combust 4.7 L methane?
How much carbon dioxide will be formed?
1
Expert's answer
2017-04-17T09:23:05-0400
CH4+2O2→CO2+2H2O
V(O_2 ):V(CH_4 )=2:1

V(O_2)=V(CH_4 )×2=4.7×2=9.4 L

V(CO_2 ):V(CH_4 )=1:1

V(CO_2)=V(CH_4 )=4.7 L

Answer:
V(O2) = 9.4 L,
V(CO2) = 4.7 L.